# Upper Approximation Bounds for Neural Oscillators

### v1

@@ 4.1 Case 1
Refer to caption(a)
Refer to caption(b)
Figure 1: Neural oscillator approximation errors $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{E_{X},2}$ versus the number $H_{{\it{\Pi}}_{i}}$ of hidden layers in ${\it{\Pi}}_{i}(\cdot)$. (a) $\tilde{\varepsilon}_{E_{X},\infty}$ versus $H_{{\it{\Pi}}_{i}}$. (b) $\tilde{\varepsilon}_{E_{X},2}$ versus $H_{{\it{\Pi}}_{i}}$.
For the five trained neural oscillators, the decays of their $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{E_{X},2}$ with respect to the increasing hidden layer numbers $H_{{{\it{\Pi}}}_{i}}$ of their MLPs ${\it{\Pi}}_{i}(\cdot)$ are shown in Figure 1. In this numerical case, each neural oscillator uses its second-order ODE to encode the input $U_{\text{e}}(t)$ and employs its MLP ${\it{\Pi}}_{i}(\cdot)$ to map the encoded features to the output function $\mathbf{y}(t)$. The results in Figure 1 illustrate that the decays of $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{X,2}$ with increasing hidden layer numbers $H_{{{\it{\Pi}}}_{i}}$ of ${\it{\Pi}}_{i}(\cdot)$ can be approximated by two curves with the same decay rate of 1/7-1/7. This value coincides with the theoretical decay rate $1/\left[p(M_{{\it{\Gamma}}}+1+1)\right]=-1/\left[1\times(5+1)+1\right]=-1/7$ in Eq. (20) of Theorem 1. Such agreement confirms the validity of the decay rate established in Theorem 1.
@@ 4.2 Case 2
$5\times 10^{4}$ independent discrete-time samples $U_{\text{e},l}(t_{i})$ of $U_{\text{e}}(t)$ are simulated. Their induced $5\times 10^{4}$ samples $X_{5,l}(t_{i})$ of the response $X_{5}(t)$ are computed. In this case, the mapping from $U_{\text{e}}(t)$ to $X_{5}(t)$ is learned using the neural oscillator with the simulated samples $U_{\text{e},l}(t_{i})$ and $X_{5,l}(t_{i})$, and the decay rate in Theorem 2 is validated.
In this case, six neural oscillators ${\it{\Pi}}_{i}\circ{\Phi}_{{\it{\Gamma}}_{i}}\left(\cdot\right)$, $i=1,2,\dots,6$, are considered. The activation functions of all MLPs in the six neural oscillators are $\sigma_{\mathrm{ReLU}}(\cdot)$. The sizes of the six one-hidden-layer MLPs ${{\it{\Pi}}}_{i}(\cdot)$ of the six neural oscillators are the same. Their widths of input, hidden, and output layers are 12, 20, and 1, respectively. The six one-hidden-layer MLPs ${{\it{\Gamma}}}_{i}(\cdot)$ of the six neural oscillators share the same input and output layer widths of 21 and 10, respectively. The hidden layer widths of ${{\it{\Gamma}}}_{1}(\cdot)$, ${{\it{\Gamma}}}_{2}(\cdot)$, ${{\it{\Gamma}}}_{3}(\cdot)$, ${{\it{\Gamma}}}_{4}(\cdot)$, ${{\it{\Gamma}}}_{5}(\cdot)$, and ${{\it{\Gamma}}}_{6}(\cdot)$ are 2, 5, 10, 20, 30, and 40, respectively.
@@
Figure 2: Neural oscillator approximation errors $\tilde{\varepsilon}_{X,\infty}$ and $\tilde{\varepsilon}_{X,2}$ versus the hidden layer width $w_{{{\it{\Gamma}}}_{i}}$ of ${\it{\Gamma}}_{i}(\cdot)$. (a) $\tilde{\varepsilon}_{X,\infty}$ versus $w_{{\it{\Gamma}}_{i}}$. (b) $\tilde{\varepsilon}_{X,2}$ versus $w_{{\it{\Gamma}}_{i}}$.
@@
For each trained neural oscillator, its $5\times 10^{4}$ prediction samples $\tilde{X}_{5,l}(t_{i})$ driven by the input samples $U_{\text{e},l}(t_{i})$ are computed. The relative approximation errors $\tilde{\varepsilon}_{X,\infty}$ and $\tilde{\varepsilon}_{X,2}$ are computed by replacing $\tilde{E}_{X_{5},l}(t_{i})$ and $E_{X_{5},l}(t_{i})$ in Eq. (36) and Eq. (37) with $\tilde{X}_{5,l}(t_{i})$ and $X_{5,l}(t_{i})$, respectively. For the six trained neural oscillators, the decays of their $\tilde{\varepsilon}_{X,\infty}$ and $\tilde{\varepsilon}_{X,2}$ with respect to the increase of hidden layer widths $w_{{\it{\Gamma}}_{i}}$ of the MLPs ${\it{\Gamma}}_{i}(\cdot)$ are shown in Figure 2. The ODE system in Eq. (33) consists of five second-order ODEs with respect to $\mathbf{X}(t)$ and five first-order ODEs with respect to $\mathbf{Z}(t)$. In this case, the each neural oscillator uses its second-order ODE to approximate the ODE system in Eq. (33) and employs its ${\it{\Pi}}_{i}(\cdot)$ as a readout function. The result in Figure 2(a) illustrates that the decay of $\tilde{\varepsilon}_{X,\infty}$ with increasing hidden layer widths $w_{{\it{\Gamma}}_{i}}$ of ${\it{\Gamma}}_{i}(\cdot)$ can be well represented by a curve with a decay rate of 0.5-0.5. This value is the same as the decay rate in Eq. (31) of Theorem 2, where the approximation error is under $L^{\infty}$-norm. The decay rate of $\tilde{\varepsilon}_{X,2}$ under $L^{2}$-norm is slightly smaller than 0.5-0.5. These results demonstrate that the approximation error decay rate in Theorem 2 is reasonable.
@@ 5 Conclusions
In this study, two upper approximation bounds of the neural oscillator consisting of a second-order ODE followed by an MLP are derived for approximating the causal and uniformly continuous operators between continuous temporal function spaces and for approximating the uniformly asymptotically incrementally stable second-order dynamical systems. The established proof method of the upper approximation bound for approximating the causal continuous operators can also be directly applied to the SS models consisting of a linear time-continuous complex RNN followed by an MLP. In this study, Theorem 1 indicates that for the causal continuous operators, the decay rate in the approximation error polynomially diminishes as the hidden-layer width of the MLP ${\it{\Gamma}}(\cdot)$ increases. Thus, the neural oscillator for approximating the causal continuous operators between continuous temporal function spaces also faces the curse of parametric complexity. Theorem 2 demonstrates that for approximating the uniformly asymptotically incrementally stable second-order dynamical systems, the neural oscillator can provide an approximation error scaling polynomially with the reciprocals of the widths of the utilized MLPs. This implies that the curse of parametric complexity can be mitigated in this situation. The approximation error decay rate with respect to the hidden layer number of the MLP ${\it{\Pi}}(\cdot)$ in Theorem 1 and that with respect to the hidden layer width of the MLP ${\it{\Gamma}}(\cdot)$ in Theorem 2 have been validated through two numerical cases.
This work only focuses on the approximation error bounds for the neural oscillator. Future research will address its generalization bounds to provide a more comprehensive theoretical understanding.

### v2

@@ 4.1 Case 1
For the five trained neural oscillators, the decays of their $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{E_{X},2}$ with respect to the increasing number of hidden layers $H_{{{\it{\Pi}}}_{i}}$ of ${\it{\Pi}}_{i}(\cdot)$ are shown in Figure 2. In this numerical case, each neural oscillator uses its second-order ODE to encode the input $U_{\text{e}}(t)$ and employs its MLP ${\it{\Pi}}_{i}(\cdot)$ to map the encoded features to the output function $\mathbf{y}(t)$. The results in Figure 2 illustrate that the decays of $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{E_{X},2}$ with the increasing number of hidden layers $H_{{{\it{\Pi}}}_{i}}$ of ${\it{\Pi}}_{i}(\cdot)$ can be approximated by two curves with the same convergence rate of 1/7-1/7. This value coincides with the theoretical convergence rate $1/\left[p(M_{{\it{\Gamma}}}+1+1)\right]=-1/\left[1\times(5+1)+1\right]=-1/7$ in Eq. (20) of Theorem 1. Such agreement validates the convergence rate established in Theorem 1.
@@ 4.2 Case 2
$5\times 10^{4}$ independent discrete-time samples $U_{\text{e},l}(t_{i})$ of $U_{\text{e}}(t)$ are simulated. Their induced $5\times 10^{4}$ samples $X_{5,l}(t_{i})$ of the response $X_{5}(t)$ are computed. In this case, the mapping from $U_{\text{e}}(t)$ to $X_{5}(t)$ is learned using the neural oscillator with the simulated samples $U_{\text{e},l}(t_{i})$ and $X_{5,l}(t_{i})$, and the convergence rate in Theorem 2 is validated.
Six neural oscillators ${\it{\Pi}}_{i}\circ{\Phi}_{{\it{\Gamma}}_{i}}\left(\cdot\right)$, $i=1,2,\dots,6$, are considered in this subsection. The activation functions of all MLPs in the six neural oscillators are $\sigma_{\mathrm{ReLU}}(\cdot)$. The sizes of the six one-hidden-layer MLPs ${{\it{\Pi}}}_{i}(\cdot)$ of the six neural oscillators are the same. Their input, hidden, and output layer widths are 12, 20, and 1, respectively. The six one-hidden-layer MLPs ${{\it{\Gamma}}}_{i}(\cdot)$ of the six neural oscillators share the same input and output layer widths of 21 and 10, respectively. The hidden-layer widths of ${{\it{\Gamma}}}_{1}(\cdot)$, ${{\it{\Gamma}}}_{2}(\cdot)$, ${{\it{\Gamma}}}_{3}(\cdot)$, ${{\it{\Gamma}}}_{4}(\cdot)$, ${{\it{\Gamma}}}_{5}(\cdot)$, and ${{\it{\Gamma}}}_{6}(\cdot)$ are 2, 5, 10, 20, 30, and 40, respectively.
@@
Figure 3: Approximation errors $\tilde{\varepsilon}_{X,\infty}$ and $\tilde{\varepsilon}_{X,2}$ versus the hidden-layer width $w_{{{\it{\Gamma}}}_{i}}$ of ${\it{\Gamma}}_{i}(\cdot)$. (a) $\tilde{\varepsilon}_{X,\infty}$ versus $w_{{\it{\Gamma}}_{i}}$. (b) $\tilde{\varepsilon}_{X,2}$ versus $w_{{\it{\Gamma}}_{i}}$.
@@
For each trained neural oscillator, its $5\times 10^{4}$ prediction samples $\tilde{X}_{5,l}(t_{i})$ driven by the input samples $U_{\text{e},l}(t_{i})$ are computed. The relative approximation errors $\tilde{\varepsilon}_{X,\infty}$ and $\tilde{\varepsilon}_{X,2}$ are computed by replacing $\tilde{E}_{X_{5},l}(t_{i})$ and $E_{X_{5},l}(t_{i})$ in Eqs. (36) and (37) with $\tilde{X}_{5,l}(t_{i})$ and $X_{5,l}(t_{i})$, respectively. For the six trained neural oscillators, the decays of their $\tilde{\varepsilon}_{X,\infty}$ and $\tilde{\varepsilon}_{X,2}$ with respect to the increase of the hidden-layer width $w_{{\it{\Gamma}}_{i}}$ of ${\it{\Gamma}}_{i}(\cdot)$ are shown in Figure 3. The ODE system in Eq. (33) consists of five second-order ODEs with respect to $\mathbf{X}(t)$ and five first-order ODEs with respect to $\mathbf{Z}(t)$. In this case, the each neural oscillator uses its second-order ODE to approximate the ODE system in Eq. (33) and employs its ${\it{\Pi}}_{i}(\cdot)$ as a readout function. The results in Figure 3(a) illustrate that the decay of $\tilde{\varepsilon}_{X,\infty}$ with the increasing hidden-layer width $w_{{\it{\Gamma}}_{i}}$ of ${\it{\Gamma}}_{i}(\cdot)$ can be well represented by a curve with a convergence rate of 0.5-0.5. This value is the same as the convergence rate in Eq. (31) of Theorem 2, where the approximation error is under the $L^{\infty}$-norm. The convergence rate of $\tilde{\varepsilon}_{X,2}$ under the $L^{2}$-norm is slightly smaller than 0.5-0.5. These results demonstrate that the convergence rate in Theorem 2 is reasonable.
4.3 Case 3
In this subsection, a zero-mean Gaussian stochastic process $U_{\text{e}}(t)$ defined over $\left[0,10\right]$ seconds is considered. Its correlation function is
${C_{\text{e}}}(t_{1},t_{2})=\exp\left\{{-200\sin^{2}{\left[0.1\left(t_{1}-t_{2}\right)\right]}}\right\}.$ (39)
The error bound $O\left[\left(\ln{M_{{\it{\Gamma}}}}\right)^{2}/{M_{\it{\Gamma}}}\right]$ in Eq. (20) of Theorem 1 is numerically validated using this Gaussian process. Specifically, 1000 samples $U_{\text{e},l}(t_{i})$ of $U_{\text{e}}(t)$ are simulated with $\Delta t=0.001$ seconds, $l=1,2,\dots,1000$. Then, a set of layer widths $M_{{\it{\Gamma}}}=\left[10,20,40,80,160,320,640,1280,2560,5000\right]^{\top}$ is considered. For each value of $M_{{\it{\Gamma}}}$, the corresponding $v\left(M_{{\it{\Gamma}}}\right)$ is calculated using Eq. (14) with $r=\lfloor\ln{M_{{\it{\Gamma}}}}\rfloor$, $c_{K}=1.1$, and $T=10$, and the corresponding 1000 approximations $\tilde{U}_{\text{e},l}(t_{i})$ are calculated by
${{\tilde{U}}_{\text{e},l}}(t_{i})={U_{\text{e},l}}(0)+\sum\limits_{n=1}^{M_{{\it{\Gamma}}}}{{\alpha_{n}}\left\{{{\mathcal{L}_{T}}{U_{\text{e},l}}(\omega_{n})+\frac{{U_{\text{e},l}}(0)}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}T)-1}\right]}\right\}\sin\left[{{\omega_{n}}(T-t_{i})-{\theta_{n}}}\right]},$ (40)
where $\omega_{n}=n\pi/\left(2T\right)$, ${\mathcal{L}_{T}}{U_{\text{e},l}}(\omega_{n})$ is the sine transform of $U_{\text{e},l}(t_{i})$ calculated by Eq. (10), $\alpha_{n}$ and $\theta_{n}$ are respectively the scaled Fourier modulus and phase of $\rho_{v\left(M_{{\it{\Gamma}}}\right)}(\tau)$ in Eq. (46) that are calculated according to Eqs. (55) and (56). The relative approximation errors $\tilde{\varepsilon}_{U,\infty}$ under the $L^{\infty}$-norm and $\tilde{\varepsilon}_{U,2}$ under the $L^{2}$-norm are computed by replacing $\tilde{E}_{X_{5},l}(t_{i})$ and $E_{X_{5},l}(t_{i})$ in Eqs. (36) and (37) with $\tilde{U}_{\text{e},l}(t_{i})$ and $U_{\text{e},l}(t_{i})$, respectively. The numerical results in Figure 4 show that the observed convergence behaviors of both $\tilde{\varepsilon}_{U,\infty}$ and $\tilde{\varepsilon}_{U,2}$ are asymptotically consistent with the theoretical error bound $O\!\left[(\ln M_{{\it{\Gamma}}})^{2}/M_{{\it{\Gamma}}}\right]$ in Theorem 1.
Refer to caption(a)
Refer to caption(b)
Figure 4: Approximation errors $\tilde{\varepsilon}_{U,\infty}$ and $\tilde{\varepsilon}_{U,2}$ versus $M_{{\it{\Gamma}}}$. (a) $\tilde{\varepsilon}_{U,\infty}$ versus $M_{{\it{\Gamma}}}$. (b) $\tilde{\varepsilon}_{U,2}$ versus $M_{{\it{\Gamma}}}$.
4.4 Case 4
In this case, a 200-degree-of-freedom Duffing system subjected to the same earthquake excitation in Case 2 is considered to illustrate the convergence properties of the neural oscillator for high-dimensional systems. Its governing differential equation is
$\mathbf{M}\mathbf{X}^{\prime\prime}(t)+\mathbf{C}\mathbf{X}^{\prime}(t)+\mathbf{F}\left[\mathbf{X}(t)\right]=-{\mathbf{M}}{U_{\text{e}}}(t),$ (41)
where ${\mathbf{X}}=\left[X_{1}(t),X_{2}(t),\dots,X_{200}(t)\right]^{\top}$, $\mathbf{X}(0)=\mathbf{X}^{\prime}(0)=\mathbf{0}$, $\mathbf{M}\in\mathbb{R}^{200\times 200}$ is a diagonal matrix $\mathbf{M}=\text{diag}(1000,1000,$ $\dots,1000)$, $\mathbf{C}\in\mathbb{R}^{200\times 200}$ is a tridiagonal matrix with $40020$ on the main diagonal and 20000-20000 on the sub- and super-diagonals, and $\mathbf{F}\left[\mathbf{X}(t)\right]=\left\{F_{1}\left[\mathbf{X}(t)\right],F_{2}\left[\mathbf{X}(t)\right],\dots,F_{200}\left[\mathbf{X}(t)\right]\right\}^{\top}$ is the restoring force vector with $F_{i}\left[\mathbf{X}(t)\right]$ being
$F_{i}\left[\mathbf{X}(t)\right]=\left\{\begin{aligned} &f_{i}\left[\mathbf{X}(t)\right]-f_{i+1}\left[\mathbf{X}(t)\right],\;\;\;\;i<200\\ &f_{i}\left[\mathbf{X}(t)\right],\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;i=200,\end{aligned}\right$. (42)
and
$f_{i}\left[\mathbf{X}(t)\right]=\left\{\begin{aligned} &10^{7}\left[X_{i}(t)-X_{i-1}(t)\right]+10^{7}\left[X_{i}(t)-X_{i-1}(t)\right]^{3},\;\;\;\;i>1\\ &10^{7}X_{i}(t)+10^{7}X_{i}^{3}(t),\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;i=1.\end{aligned}\right$. (43)
The ODE system in Eq. (41) is solved numerically using a fourth-order Runge-Kutta method (Griffiths and Higham, 2010) with a time discretization of $\Delta t=0.005$ seconds over the time interval $[0,20]$.
In this case, seven neural oscillators ${\it{\Pi}}_{i}\circ{\Phi}_{{\it{\Gamma}}_{i}}\left(\cdot\right)$, $i=1,2,\dots,7$, are utilized to learn the mapping from $U_{\text{e}}(t)$ to $X_{200}(t)$. The activation functions of all MLPs in the seven neural oscillators are $\sigma_{\mathrm{ReLU}}(\cdot)$. The seven one-hidden-layer MLPs ${{\it{\Gamma}}}_{i}(\cdot)$ of the seven neural oscillators share the same input and output layer widths of 41 and 20, respectively. The seven one-hidden-layer MLPs ${{\it{\Pi}}}_{i}(\cdot)$ share the same input and output layer widths of 22 and 1, respectively. For each neural oscillator ${\it{\Pi}}_{i}\circ{\Phi}_{{\it{\Gamma}}_{i}}\left(\cdot\right)$, ${{\it{\Gamma}}}_{i}(\cdot)$ and ${{\it{\Pi}}}_{i}(\cdot)$ share the same hidden-layer width $w_{i}$. The values of $w_{i}$ corresponding to $i=1,2,\dots,7$ are 2, 5, 10, 20, 30, 40, and 60, respectively.
The $10^{4}$ samples $U_{\text{e},l}(t_{i})$ of input $U_{\text{e}}(t)$ are simulated and arranged in a descending order with respect to their standard deviations. Then the $1^{\text{st}}$, $6^{\text{th}}$, $11^{\text{st}}$$\dots,$ $9996^{\text{th}}$ input samples and their induced samples $X_{200,l}(t_{i})$ of $X_{200}(t)$ (total 2000 input and response sample pairs) are selected for training. Among these 2000 sample pairs, the $5^{\text{th}}$, $10^{\text{th}}$, $15^{\text{th}}$$\dots,$ $2000^{\text{th}}$ pairs (total 400 pairs) are selected as a validation dataset, and the remaining 1600 sample pairs constitute a training dataset. Each of the seven neural oscillators is trained using the training approach outlined in E with these 2000 sample pairs. The mean squared loss function $\ell_{2}$ in Eq. (122) is used. The training phase comprises 10000 epochs, with a batch size of 1600 for each epoch. The initial learn rate is 0.02. The learning rate drop period is 100 epochs. The learning rate drop factor is set to 0.96. A gradient norm clipping strategy (Pascanu et al., 2013) with a threshold of 1 is applied to all MLPs. The parameters of all MLPs are randomly initialized. For each neural oscillator, the neural network after the last update is selected as the final model.
Refer to caption(a)
Refer to caption(b)
Figure 5: Approximation errors $\tilde{\varepsilon}_{X_{200},\infty}$ and $\tilde{\varepsilon}_{X_{200},2}$ versus the hidden-layer widths $w_{i}$ of ${\it{\Gamma}}_{i}(\cdot)$ and ${\it{\Pi}}_{i}(\cdot)$. (a) $\tilde{\varepsilon}_{X_{200},\infty}$ versus $w_{i}$. (b) $\tilde{\varepsilon}_{X_{200},2}$ versus $w_{i}$.
For each trained neural oscillator, its $10^{4}$ prediction samples $\tilde{X}_{200,l}(t_{i})$ driven by the input samples $U_{\text{e},l}(t_{i})$ are computed. The relative approximation errors $\tilde{\varepsilon}_{X_{200},\infty}$ and $\tilde{\varepsilon}_{X_{200},2}$ are computed by replacing $\tilde{E}_{X_{5},l}(t_{i})$ and $E_{X_{5},l}(t_{i})$ in Eqs. (36) and (37) with $\tilde{X}_{200,l}(t_{i})$ and $X_{200,l}(t_{i})$, respectively. For the seven trained neural oscillators, the decays of their $\tilde{\varepsilon}_{X_{200},\infty}$ and $\tilde{\varepsilon}_{X_{200},2}$ with respect to the increase of the hidden-layer widths $w_{i}$ of ${\it{\Gamma}}_{i}(\cdot)$ and ${\it{\Pi}}_{i}(\cdot)$ are shown in Figure 5. The results illustrate that the decays of $\tilde{\varepsilon}_{X_{200},\infty}$ and $\tilde{\varepsilon}_{X_{200},2}$ with respect to $w_{i}$ can be approximately represented by two curves with the same convergence rate of 0.5-0.5, that is consistent with the result in Theorem 2. Furthermore, the relative errors $\tilde{\varepsilon}_{X_{200},\infty}<0.1$ and $\tilde{\varepsilon}_{X_{200},2}<0.05$ are achieved by the model with $w_{i}=30$. The probability distribution function (PDF) and cumulative distribution function (CDF) of the extreme value $E_{\text{max}}=\max_{t\in\left[0,20\right]}\left|X_{200}(t)\right|$ estimated from the $10^{4}$ samples $\tilde{X}_{200,l}(t_{i})$ by the the model with $w_{i}=30$ are consistent with the results from the samples $X_{200,l}(t_{i})$, as shown in Figure 6. These results suggest that, since the number of the dominant modes governing the response of the high-dimensional system in Eq. (41) is limited, a neural oscillator with a moderate network size can effectively approximate the system. From Figure 5, the comparison between the errors from the model with $w_{i}=30$ and those from the model with $w_{i}=60$ indicates that, given the convergence rate of 0.5-0.5, the expressive power of the neural oscillator improves sublinearly with respect to the hidden-layer width $w_{i}$. As a result, the marginal benefit of increasing model size gradually diminishes. Therefore, excessively large networks may not yield proportional performance improvements, and their practical advantage may be limited relative to their computational cost.
Refer to caption(a)
Refer to caption(b)
Figure 6: PDF and CDF of $E_{\text{max}}$. (a) PDF of $E_{\text{max}}$. (b) CDF of $E_{\text{max}}$.
@@ 5 Conclusions
In this study, two upper approximation bounds of the neural oscillator consisting of a second-order ODE followed by an MLP are derived for approximating the causal and uniformly continuous operators between continuous temporal function spaces and for approximating the uniformly asymptotically incrementally stable second-order dynamical systems. The established proof method of the upper approximation bound for approximating the causal continuous operators can also be directly applied to the SS models consisting of a linear time-continuous complex RNN followed by an MLP. Theorem 1 indicates that for the causal continuous operators, the convergence rate in the approximation error polynomially diminishes as the hidden-layer width of the MLP ${\it{\Gamma}}(\cdot)$ increases. Thus, the neural oscillator for approximating the causal continuous operators between continuous temporal function spaces also faces the curse of parametric complexity. Theorem 2 demonstrates that for approximating the uniformly asymptotically incrementally stable second-order dynamical systems, the neural oscillator can provide an approximation error scaling polynomially with the reciprocals of the widths of the utilized MLPs. This implies that the curse of parametric complexity can be overcome in this situation. Four numerical cases are presented to validate the convergence rates of the derived approximation error bounds and to illustrate the convergence properties of the neural oscillator. The results show that the convergence rate with respect to the number of hidden layers of the MLP ${\it{\Pi}}(\cdot)$ in Theorem 1 and those with respect to the hidden-layer widths of the MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ in Theorems 1 and 2 are consistent with the numerical results. A neural oscillator with a moderate network size can effectively approximate high-dimensional systems characterized by a small number of dominant modes.
@@
Data and Code Availability
All code used for the numerical experiments in this study is available at: https://github.com/ZifengH22/Upper-Approximation-Bounds-for-Neural-Oscillators. All data presented in this study were generated using this code.